An analysis of the monthly allowances received by a sample of five engineers


from a telco company has been conducted recently. The mean and median of


the allowances is RM7000. A unimodal among the observations is RM12,000.


Find the standard deviation if the allowances paid to each engineer were in full


thousands

1
Expert's answer
2022-02-16T16:43:05-0500

Mean "\mu=7000"

x=12000

N=5


So,

"1000=\dfrac{x-\mu}{\frac{\sigma}{\sqrt{n}}}\\\implies \dfrac{\sigma}{\sqrt{5}}=\dfrac{12000-7000}{1000}\\\implies \sigma=5\times \sqrt{5}=11.18"